OpenStudy (anonymous):

what algebraic equation represents "three times the differece of a number,x, and nine equals fifteen

OpenStudy (anonymous):

3(x - 9) = 15

Parth (parthkohli):

3(x-9) = 15

OpenStudy (anonymous):

3(x - 9) = 15

OpenStudy (anonymous):

the difference of x and 9 is 9-x if x is less than 9, and x-9 if x is bigger than 9 so we want 3 times the difference 3(9-x) and this equals 15 3(9-x) = 15 and if x is bigger than 9 we have 3(x-9) = 15

OpenStudy (anonymous):

3 (x-9) = 15

OpenStudy (anonymous):

@eigenschmeigen is correct

OpenStudy (anonymous):

3 |x-9|=15

OpenStudy (anonymous):

so is @Thomas9 i wasn't sure if the asker had come across the absolute value function

OpenStudy (anonymous):

are you happy with these explanations sps?